OpenStudy (ccboo24):

What is the vertex of the parabola defined by the equation y = x2 + 4x + 4?

Nnesha (nnesha):

the formula to find x-coordinate of the vertex is \[\rm x=\frac{-b}{2a}\] a,b,c values are \[\rm Ax^2+Bx+C=0\] where a=leading coefficient b= coefficient of x term c=constant find x and then substitute this x value into the original quadratic equation to find y coordinate

OpenStudy (ccboo24):

so would it be 2,0?

OpenStudy (michele_laino):

hint: \(a=1\) and \(b=4\) so: \[{x_V} = \frac{{ - b}}{{2a}} = \frac{{ - 4}}{{2 \cdot 1}} = ...?\]

OpenStudy (michele_laino):

wherein \(x_V\) stands for x-coordinate of the vertex

OpenStudy (ccboo24):

1,2?

Nnesha (nnesha):

how did you get that ??

Nnesha (nnesha):

i know there are 4 options: first one is ) 2,0 2nd) 1,2 3rd) idk 4th) idk ;) don't guess! c;

OpenStudy (ccboo24):

there's 2,0 1,2 -2,0 4,2 0,0

Nnesha (nnesha):

okay please plugin b and a value into the equation just like in 3rd comment on this post and then simplify the fraction

OpenStudy (ccboo24):

4,2

Nnesha (nnesha):

how did you get that ? what's -4/2 = ?